OpenStudy (anonymous):

ln 3 − 2(ln 4 + ln 8)

OpenStudy (anonymous):

lna + lnb = lna*b and lna- lnb = lna/b so ln4+ln8 = ln4*8 = ln32 ln3 -2ln32 = ln3-ln32^2 (another rule alnb = lnb^a) so = ln3/(ln32^2)